OpenStudy (muhammad_nauman_umair):

What will be \[\int\limits {(4x*y^2+6*y)/(-x^2y^2-2*x*y)} dx\]

OpenStudy (amistre64):

is that y(x)? or is it independant as well?

OpenStudy (amistre64):

4x+6 ------- seems to be whats important if its a partial -x^2-2x

OpenStudy (amistre64):

i dont spose we can decomp that .....

sam (.sam.):

I think you need to use partial fractions

OpenStudy (muhammad_nauman_umair):

In it both x and y are present

OpenStudy (anonymous):

make variable change xy = z, and its easy

OpenStudy (muhammad_nauman_umair):

SAM how can we use partial fraction in it

OpenStudy (anonymous):

xy =z, z' = y, 2zz' =2xy2

OpenStudy (anonymous):

u get Ln(z2 +z)dz, and come back to xy

OpenStudy (anonymous):

got it?

OpenStudy (muhammad_nauman_umair):

wait im tryin myko

OpenStudy (anonymous):

kk

sam (.sam.):

from,cancel common terms in the num and denom \[\begin{array}{l} \text{ }\frac{4 x y^2+6 y}{-x^2 y^2-2 x y}\text{ } \\ \text{} \int\limits -\frac{2 (2 x y+3)}{x (x y+2)} \, dx \\\end{array}\] then use partial fractions: you'll get \[-2\int\limits \left(\frac{y}{2 (x y+2)}+\frac{3}{2 x}\right) \, dx\] then integrate using u-sub check if im wrong

OpenStudy (anonymous):

\[\int\limits[2z+1/z ^{2}+z]dz = Ln (z ^{2}+z) + C\]

OpenStudy (anonymous):

sry, forgot - sign

sam (.sam.):

@Zarkon What's your opinion?

OpenStudy (anonymous):

in front of integral

OpenStudy (zarkon):

you are correct Sam

OpenStudy (anonymous):

to complicated

sam (.sam.):

\[Ans: -\ln (x y+2)-3 \ln (x)\]

OpenStudy (anonymous):

\[\int\limits(2z+2/z ^{2}+2z)dz +\int\limits(4/z ^{2}+2z) dz = Ln (z ^{2}+2z)+ 4Ln(z+2)\]